OpenStudy (anonymous):

solve each system by substitution. 4x+4y=0 y=2x+6

OpenStudy (richyw):

I explained basically this exact question to you earlier!

OpenStudy (anonymous):

whats x and y?

OpenStudy (anonymous):

4x + 4(2x-6) = 0 Just plug in 2x+6 everywhere you see y

OpenStudy (anonymous):

ya there bud

OpenStudy (anonymous):

4x + 8x + 24 = 0 Combine like terms: 12x + 24 = 0 Subtract 24 from both sides: 12x = -24 Divide each side by 12: x = -2 Then plug it in to solve for y: 4(-2) + 4y = 0 -8 + 4y = 0 Add 8 to both sides: 4y = 8 Divide each side by 4: y = 2 So the answer is (-2, 2)

OpenStudy (precal):